OpenStudy (anonymous):

Find the limit of [1+(2/x^2)]^2x as x approaches infinity

OpenStudy (anonymous):

Compare this to the limit that yields e

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(1+\frac{2}{x^2}\right)^{2x}\\ \large e^{\ln\left[\lim_{x\to\infty}\left(1+\frac{2}{x^2}\right)^{2x}\right]}\\ \large e^{\lim_{x\to\infty}\ln\left(1+\frac{2}{x^2}\right)^{2x}}\\ \large e^{\lim_{x\to\infty}2x\ln\left(1+\frac{2}{x^2}\right)}\] Focusing on just the exponent for a moment: \[\lim_{x\to\infty}2x\ln\left(1+\frac{2}{x^2}\right)\\ 2\lim_{x\to\infty}\frac{\ln\left(1+\frac{2}{x^2}\right)}{\frac{1}{x}}=\frac{0}{0}\] Applying L'Hopital's rule, you get \[2\huge \lim_{x\to\infty}\frac{ \frac{-\frac{4}{x^3}}{1+\frac{2}{x^2}} }{-\frac{1}{x^2}}\\ 2\large \lim_{x\to\infty}\frac{ \frac{4}{x} }{ 1+\frac{2}{x^2}}=\frac{0}{1}=0\\ \]So, you have \[\large e^{\displaystyle \lim_{x\to\infty}\cdots}=e^0=1\]